OpenStudy (lexilove13):

2^x=1/8 Can someone please explain to me how to solve this?

OpenStudy (retireed):

I did it graphically, and got 2 to a negative exponent I am not ready to post yet. I think you need to take the log of both sides, but I need to try it to be sure.

Nnesha (nnesha):

^ ^ yes take log both sides that will help to bring down the variable \[ y^x \rightarrow x \ln y\]

OpenStudy (lexilove13):

@Nnesha Can you explain like i am in 2nd grade? I am so lost and need to learn this for a test tomorrow

OpenStudy (retireed):

log 2^x = log (0.125) x = log(0.125) / log(2)

OpenStudy (lexilove13):

OHHHHHHH ok! Now i get it! Thank you @retirEEd

OpenStudy (retireed):

x will equal -3 and I glad I could remember how to do it!